# Variance-Reducing Couplings for Random Features: Perspectives from Optimal Transport

### v1

@@ B.2 Do better kernel approximations give better predictions?
Figure 5: Results on a single train-test split of the power dataset, showing kernel approximation RMSE, KL divergence to the true predictive posterior, test RMSE of predictions and KL divergence to the true prior. We have successfully reduced the variance of the kernel approximation, but this may not help downstream metrics. Standard deviation estimates are shaded but too small to easily see.
@@ Appendix C Graph random features
In this appendix, we provide a self-contained introduction to graph random features (GRFs) and previously-proposed techniques to reduce their variance. We also explain the motivations behind the series of approximations used in Sec. 4.1, and where possible provide empirical evidence that these approximations for tractability and efficiency do not substantially degrade the performance of the learned $\sigma$-coupling.
@@ C.1 Constructing graph random features
For the reader’s convenience, we begin by providing a brief self-contained introduction to GRFs.
@@
where $\bm{\omega}_{k}^{(i)}$ is the $k$th walk (out of a total of $m$) simulated out of the starting node $v_{i}$. We remind the reader that by a ‘walk’ we mean a sequence of neighbouring graph nodes. The projection function $\psi(\cdot):\Omega\to\mathbb{R}^{N}$ maps from then set of graph random walks $\Omega\coloneqq\left\{(v_{i})_{i=1}^{l}\,|\,v_{i}\in\mathcal{N},(v_{i},v_{i+1})\in\mathcal{E},l\in\mathbb{N}\right\}$ to a sparse $N$-dimensional feature vector. It is computed as follows:
@@ C.2 Antithetic termination
Consider a random walker that terminates with probability $p$ at every timestep. In practice, this can be achieved by drawing a ‘termination random variable’ $t$ from a uniform distribution on $[0,1]$, $t\sim\mathcal{U}([0,1])$, and ending the walk if $t<p$. Now consider two such walkers with corresponding termination random variables $t_{1},t_{2}$. If they are independent, $t_{1}$ and $t_{2}$ are drawn independently. Antithetic termination instead proposes to induce a nontrivial joint distribution by first drawing $t_{1}\sim\mathcal{U}([0,1])$ as usual, and then setting $t_{2}=\mod_{1}(t_{1}+\frac{1}{2})$. $t_{2}$ still follows a uniform marginal distribution so we preserve unbiasedness, but this coupling forces walkers to terminate at different timesteps (if $p<0.5$) which the authors prove reduces the estimator variance under certain assumptions. This is one example of a possible hand-crafted coupling between random walk lengths that improves estimator concentration.
As briefly noted in the main text, it is also possible to improve the accuracy of estimators using graph random walks by coupling walker directions [Reid et al., 2024a, Luo, 2019]. This approach is complementary to our own and can be combined with it. Formulating coupled walker directions in terms of OT is an interesting and involved technical challenge; we defer its study to future work.
@@
[…]
grfs_master_schematic
Figure 6: The left schematic shows a permutation density $p_{\sigma}$ for the permutation $\sigma=$24351, corresponding to a measure $\pi^{(\sigma)}$ on $[0,1]^{2}$. Both coordinates are pushed forward with $F_{\eta_{\textrm{G}}}^{-1}$ to get the coupling $\mu^{(\sigma)}$. Then the permutation can be interpreted as a matching between the quantiles of the geometric distribution over walk lengths, sketched in the central schematic. We use this to couple the lengths of pairs of walkers, shown on the right. Taking these walkers as inputs to the projection functions $\psi(\cdot)$ (Eq. 10), we get $\sigma$-coupled GRFs.
In this appendix, we discuss the steps to formulate the OT matching problem in Sec. 4.1. We will especially focus on the approximations needed to make the objective in Eq. 13 tractable, describing their motivation and where possible investigating how they modify the original objective. For reference, Fig. 6 gives an overview of the coupling mechanism.
@@
$\mu^{*}=\arg\inf_{\mu\in\Lambda_{2}(\eta_{\textrm{G}})}\mathbb{E}_{(l_{1}^{(i)},l_{2}^{(i)})\sim\mu,(l_{1}^{(j)},l_{2}^{(j)})\sim\mu}\left(\mathbb{E}_{\textrm{dirs}}\left[\left(\psi(\bm{\omega}_{1}^{(i)})+\psi(\bm{\omega}_{2}^{(i)})\right)^{\top}\left(\psi(\bm{\omega}_{1}^{(j)})+\psi(\bm{\omega}_{2}^{(j)})\right)\right]^{2}\right),$ (43)
@@
This computes the average feature projection over all walks of a given length $l^{(i)}$ out of node $v_{i}$. It is straightforward to estimate by simulating a finite number of random walks. If we move the expectation over walker directions inside the square bracked, the (now approximate) OT problem becomes
$\mu^{*}=\arg\inf_{\mu\in\Lambda_{2}(\eta_{\textrm{G}})}\mathbb{E}_{(l_{1}^{(i)},l_{2}^{(i)})\sim\mu,(l_{1}^{(j)},l_{2}^{(j)})\sim\mu}\left[\left(\widetilde{\psi}(l_{1}^{(i)})+\widetilde{\psi}(l_{2}^{(i)})\right)^{\top}\left(\widetilde{\psi}(l_{1}^{(j)})+\widetilde{\psi}(l_{2}^{(j)})\right)\right]^{2}.$ (45)
@@
Using $\sigma$-couplings. Unfortunately, Eq. 45 remains intractable. To make progress, we must restrict our search-space of joint distributions from $\Lambda_{2}(\eta_{\textrm{G}})$ – the set of all joint distributions with geometrically-distributed marginals – to a tractable subclass. Taking inspiration from approaches in numerical OT, we have considered the class of permutation densities defined in Eq. 14, transformed coordinate-wise by pushing forward with the appropriate inverse CDF. This may initially seem an arbitrary choice, but it is motivated by three important observations. First, we have seen that this class of joint distributions frames the minimisation as a matching problem which, under further simplifying assumptions, we can efficiently solve using techniques from linear programming. Second, it is straightforward to sample from so the learned coupling will be practically useful (see comments in Sec. 2). Third, it is well-motivated by results in discrete OT; see App. D for concrete theoretical guarantees. Once again, we have seen that in practice, even at modest permutation order $n$, this class contains couplings that can substantially outperform both i.i.d. walkers and antithetic termination [Reid et al., 2024c].
We now replace $\Lambda_{2}(\eta_{\textrm{G}})$ in Eq. 45 by the set $\{\mu^{(\sigma)}\}_{\sigma\in S_{n}}$, with $S_{n}$ the set of permutations $[\![n]\!]\to[\![n]\!]$. Denote by $\widehat{\psi}(\cdot):[\![n]\!]\to\mathbb{R}^{N}$ the function
$\widehat{\psi}(q^{(i)})\coloneqq\mathbb{E}_{u\sim\mathcal{U}([\frac{q-1}{n},\frac{q}{n}])}\left(\widetilde{\psi}(F^{-1}_{\eta_{\textrm{G}}}(u)^{(i)})\right).$ (46)
@@
$\sigma^{*}=\arg\inf_{\sigma}\sum_{q_{1}\in[\![n]\!]}\sum_{q_{2}\in[\![n]\!]}\left[\left(\widehat{\psi}(q_{1}^{(i)})+\widehat{\psi}(\sigma(q_{1})^{(i)})\right)^{\top}\left(\widehat{\psi}(q_{2}^{(j)})+\widehat{\psi}(\sigma(q_{2})^{(j)})\right)\right]^{2}$ (47)
@@
Solving the matching problem. Though we can now efficiently estimate the cost function with MC, it is is still difficult to solve the matching problem in Eq. 47 efficiently because the cost function weights are quadratic. Minimum weights quadratic bipartite matching problems are generally NP-hard, but here the problem has extra symmetry that makes it simpler.
In App. C.3.1, we discuss possible approaches to directly solve the matching problem in Eq. 47. In particular, we prove that it can be solved with high probability at a time complexity independent of the number of nodes $N$, and give a polynomial time algorithm for the special case of diagonal kernel entries $k(v_{i},v_{i})$. Notwithstanding this progress, as a simpler first recourse we can just make a further approximation by restricting the sum to $q_{1}=q_{2}$ diagonal terms. This is the approximation used in the main text. Doing so, we arrive at
$\sigma^{*}=\arg\inf_{\sigma}\sum_{q\in[\![n]\!]}\left[\left(\widehat{\psi}(q^{(i)})+\widehat{\psi}(\sigma(q)^{(i)})\right)^{\top}\left(\widehat{\psi}(q^{(j)})+\widehat{\psi}(\sigma(q)^{(j)})\right)\right]^{2}.$ (48)
@@
C.3.1 Addendum: towards an analytic solution to Eq. 47
To assuage any possible dissatisfaction with the $q_{1}=q_{2}$ approximation in Eq. 48, in this section we discuss progress towards solving the optimisation problem in Eq. 47 exactly. In particular, we prove that it can be solved with high probability and:
@@
    at a time complexity independent from the dimensionality of vectors $\{\widehat{\psi}(q^{(i,j)})\}_{q=1}^{n}$ (at the cost of the one-time pre-processing), or
@@
First, for clarity of exposition, we will introduce more convenient notation and important definitions. For a given $\epsilon>0$ and a vector $\mathbf{v}\in\mathbb{R}^{d}$, define the set $N_{\epsilon}(\mathbf{v})=\{\mathbf{w}\in\mathbb{R}^{d}:|\alpha_{\mathbf{v},\mathbf{w}}-\frac{\pi}{2}|\leq\epsilon\}$, where $\alpha_{\mathbf{v},\mathbf{w}}$ is an angle between $\mathbf{v}$ and $\mathbf{w}$. Then define the property of $\epsilon$-separation as follows.

### v2

@@ B.2 PNC RFFs for Gaussian processes
Figure 6: Results on a single train-test split of the power dataset, showing kernel approximation RMSE, KL divergence to the true predictive posterior, test RMSE of predictions and KL divergence to the true prior. We have successfully reduced the variance of the kernel approximation, but this may not help downstream metrics. Standard errors are shaded but too small to easily see.
B.3 Norm-coupled RLFs in Performers
In Sec. 3.3, we used random Laplace features for attention approximation in Performers [Choromanski et al., 2020], a type of efficient transformer that estimates softmax using a low rank decomposition. Here, we discuss the results in greater detail.
Derivation of Eq. 8. Let $X_{ij}$ denote the random variable $\widehat{k}(\bm{x}_{i},\bm{x}_{j})$, which is an unbiased estimate of $\exp(\bm{x}_{i}^{\top}\bm{x}_{j})$ constructed using features $\{\exp({\|\bm{x}_{i}\|^{2}}/{2})\phi_{\textrm{RLF}}(\bm{x}_{i})\}_{i=1}^{N}$. Unbiasedness follows trivially from the fact that RLFs give an unbiased estimate of the Gaussian kernel. Normalising by the sum of attention scores, let us define $\widehat{a}_{ij}\coloneqq X_{ij}/\sum_{j}X_{ij}$. Let $X_{ij}=\mu_{ij}+\delta_{ij}$ with $\mu_{ij}\coloneqq\mathbb{E}(X_{ij})$ and $\mathbb{E}(\delta_{ij})=0$. Then
[…]
where we defined $\bar{\mu}_{i}\coloneqq\frac{1}{N}\sum_{j}\mu_{ij}$, the average groundtruth attention score across tokens. Since we are targeting $\frac{\mu_{ij}}{N\bar{\mu_{i}}}$,
$\textrm{MSE}(\widehat{a}_{ij})=\frac{1}{N^{2}\bar{\mu_{i}}^{2}}\left(\delta_{ij}^{2}-\frac{4\mu_{ij}\delta_{ij}}{N\bar{\mu}_{i}}\sum_{k}\delta_{ik}+\frac{3\mu_{ij}^{2}}{N^{2}\bar{\mu}_{i}^{2}}\sum_{kl}\delta_{ik}\delta_{il}\right)+\mathcal{O}\left(\frac{1}{N^{3}}\right).$ (45)
As in the main text, denote $\textrm{MSE}(\widehat{a}_{i})\coloneqq\frac{1}{N}\sum_{j}\textrm{MSE}(\widehat{a}_{ij})$, the average mean squared error over the tokens to which $i$ attends. To better see the intuitive behaviour, suppose also that $\mu_{ij}=\bar{\mu}_{i}\hskip 1.99997pt\forall\hskip 1.99997ptj$. Then
$\end{multlined}\textrm{MSE}(\widehat{a}_{i})=\frac{1}{N^{2}\bar{\mu_{i}}^{2}}\left(\frac{1}{N}\sum_{j}\delta_{ij}^{2}-\frac{1}{N^{2}}\sum_{jk}\delta_{ij}\delta_{ik}\right)+\mathcal{O}\left(\frac{1}{N^{3}}\right)\\ =\frac{1}{N^{2}\bar{\mu}_{i}^{2}}\left(\frac{1}{N}\sum_{j}\textrm{Var}(\widehat{k}(\bm{x}_{i},\bm{x}_{j}))-\frac{1}{N^{2}}\sum_{j,k}\textrm{Cov}(\widehat{k}(\bm{x}_{i},\bm{x}_{j}),\widehat{k}(\bm{x}_{i},\bm{x}_{k}))\right)+\mathcal{O}\left(\frac{1}{N^{3}}\right),$ (46)
as reported in Eq. 8.
Positive monotone coupling. In contrast to Eq. 7, we say that random variables $(\omega_{1},\omega_{2})$ are postive monotone-coupled if $\omega_{1}=\omega_{2}$ almost surely. This is a valid transport plan, the identity. Clearly, all $m$ frequencies in an ensemble can be simultaneously positive monotone-coupled by making them equal. It is intuitive that this should maximise the kernel estimator variance since we sample only one frequency lengthscale. The proof is a trivial extension of the arguments made for Thm. 3.2 in App. A.2, simply taking $c_{\textrm{RLF}}\to-c_{\textrm{RLF}}$ so that the support of the OT plan swaps. We omit it for brevity. It is also intuitive that this will increase the covariance between kernel estimates.
Variance, covariance, and attention approximation error. Modifying the norm coupling between frequencies changes both $\textrm{Var}(\widehat{k}(\bm{x}_{i},\bm{x}_{j}))$ and $\textrm{Cov}(\widehat{k}(\bm{x}_{i},\bm{x}_{j}))$, so $\textrm{MSE}(\widehat{a}_{i})$ can change unpredictably. Fig. 7 shows the results for randomly $N=16$ synthetic, normally distributed $16$-dimensional keys, $\bm{x}\sim\mathcal{N}(0,\frac{1}{\sqrt{d}}\mathbf{I}_{d}$). Strikingly, maximising the kernel estimator variance with a positive monotone (PM) coupling ends up improving the quality of attention estimation since it also also increases the covariance between the unnormalised scores. This counterintuitive finding shows highlights the limitations of variance reduction as a paradigm.
Figure 7: Kernel variance $\textrm{Var}(\widehat{k}(\bm{x}_{i},\bm{x}_{j}))$, kernel covariance $\textrm{Cov}(\widehat{k}(\bm{x}_{i},\bm{x}_{j}))$ and attention mean squared error $\textrm{MSE}(a_{i})$ plotted against number of frequencies for different coupling schemes. As intended, PNC decreases the kernel estimator variance, but it also decreases the kernel estimator covariance so the attention MSE remains essentially unchanged. On the other hand, positive monotone coupling (PM) drastically increases both kernel variance (which it maximises) and kernel covariance. The latter offsets the former (Eq. 46), so that attention MSE actually drops: a remarkable, counterintuitive finding.
Performer experimental details. To obtain the results in Table 2, we train a Performer-ViT [Dosovitskiy et al., 2020, Choromanski et al., 2020] on the ImageNet (1M) dataset [Deng et al., 2009]. We estimate attention using $m=128$ random features that are orthogonal with norms that are: (i) i.i.d. (ii) PNC (Def. 3.3), or (iii) positive monotone-coupled (i.e. equal among blocks of size $d$). We use a transformer with $12$ layers and heads, with hidden size $768$ and MLP dimension $3072$. We take $16\times 16$ patches, and train with the Adam optimiser for $90$ epochs with a compound learning rate ($10^{4}$ steps linear warmup, constant, then cosine decay, with base LR $3\times 10^{-3}$ and final LR $1\times 10^{-5}$). The batch size is $4096$. To get the standard errors, we average between $34$ and $48$ seeds per coupling. We see that the lower attention MSE unlocked by PM coupling improves predictive performance over the baseline.
@@ Appendix C Graph random features
In this appendix, we provide a self-contained introduction to graph random features (GRFs) and previously proposed techniques to reduce their variance. We also explain the motivations behind the series of approximations used in Sec. 4.1, and where possible provide empirical evidence that these approximations for tractability and efficiency do not substantially degrade the performance of the learned $\sigma$-coupling.
@@ C.1 Constructing graph random features
For the reader’s convenience, we begin by providing a brief introduction to GRFs.
@@
where $\bm{\omega}_{k}^{(i)}$ is the $k$th walk (out of a total of $m$) simulated out of the starting node $v_{i}$. We remind the reader that by a ‘walk’ we mean a sequence of neighbouring graph nodes. The projection function $\psi(\cdot):\Omega\to\mathbb{R}^{N}$ maps from the set of graph random walks $\Omega\coloneqq\left\{(v_{i})_{i=1}^{l}\,|\,v_{i}\in\mathcal{N},(v_{i},v_{i+1})\in\mathcal{E},l\in\mathbb{N}\right\}$ to a sparse $N$-dimensional feature vector. It is computed as follows:
@@ C.2 Antithetic termination
Consider a random walker that terminates with probability $p$ at every timestep. In practice, this can be achieved by drawing a ‘termination random variable’ $t$ from a uniform distribution on $[0,1]$, $t\sim\mathcal{U}([0,1])$, and ending the walk if $t<p$. Now consider two such walkers with corresponding termination random variables $t_{1},t_{2}$. If they are independent, $t_{1}$ and $t_{2}$ are drawn independently. Antithetic termination instead proposes to induce a nontrivial joint distribution by first drawing $t_{1}\sim\mathcal{U}([0,1])$ as usual, and then setting $t_{2}=\mod_{1}(t_{1}+\frac{1}{2})$. $t_{2}$ still follows a uniform marginal distribution so we preserve unbiasedness, but this coupling forces walkers to terminate at different timesteps (if $p_{\textrm{halt}}<0.5$) which the authors prove reduces the estimator variance under certain assumptions. This is one example of a possible hand-crafted coupling between random walk lengths that improves estimator concentration.
It is also possible to improve the accuracy of estimators using graph random walks by coupling walker directions [Reid et al., 2024a, Luo, 2019]. This approach is complementary to our own and can be combined with it. Formulating coupled walker directions in terms of OT is an interesting and involved technical challenge; we defer its study to future work.
@@
In this appendix, we discuss the steps to formulate the OT matching problem in Sec. 4.1. We will especially focus on the approximations needed to make the objective in Eq. 12 tractable, describing their motivation and where possible investigating how they modify the original objective.
@@
minimise $\end{multlined}\textrm{minimise }\mathbb{E}_{(l_{1}^{(i)},l_{2}^{(i)})\sim\mu,(l_{1}^{(j)},l_{2}^{(j)})\sim\mu}\left(\mathbb{E}_{\textrm{dirs}}\left[\left(\psi(\bm{\omega}_{1}^{(i)})+\psi(\bm{\omega}_{2}^{(i)})\right)^{\top}\left(\psi(\bm{\omega}_{1}^{(j)})+\psi(\bm{\omega}_{2}^{(j)})\right)\right]^{2}\right)\\ \textrm{for }\mu\in\Lambda_{2}(\eta_{\textrm{G}}),$ (51)
@@
This computes the average feature projection over all walks of a given length $l^{(i)}$ out of node $v_{i}$. It is straightforward to estimate by simulating a finite number of random walks. If we move the expectation over walker directions inside the square bracket, the (now approximate) OT problem becomes
minimise $\textrm{minimise }\mathbb{E}_{(l_{1}^{(i)},l_{2}^{(i)})\sim\mu,(l_{1}^{(j)},l_{2}^{(j)})\sim\mu}\left[\left(\widetilde{\psi}(l_{1}^{(i)})+\widetilde{\psi}(l_{2}^{(i)})\right)^{\top}\left(\widetilde{\psi}(l_{1}^{(j)})+\widetilde{\psi}(l_{2}^{(j)})\right)\right]^{2}\textrm{ for }\mu\in\Lambda_{2}(\eta_{\textrm{G}}).$ (53)
@@
Using $\sigma$-couplings. Unfortunately, Eq. 53 remains intractable. To make progress, we must restrict our search-space of joint distributions from $\Lambda_{2}(\eta_{\textrm{G}})$ – the set of all joint distributions with geometrically-distributed marginals – to a tractable subclass. Taking inspiration from approaches in numerical OT, we have considered the class of permutation densities defined by $p_{\sigma}(x,y)\coloneqq n1_{\sigma(\lceil nx\rceil)=\lceil ny\rceil}$, with $\sigma$ a permutation of order $n$ (that is, a bijection $\sigma:[\![n]\!]\to[\![n]\!]$), transformed coordinate-wise by pushing forward with the appropriate inverse CDF. This may initially seem an arbitrary choice, but it is motivated by three important observations. First, we have seen that this class of joint distributions frames the minimisation as a matching problem which, under further simplifying assumptions, we can efficiently solve using techniques from linear programming. Second, it is straightforward to sample from so the learned coupling will be practically useful. Third, it is well-motivated by results in discrete OT; see App. D for concrete theoretical guarantees. Once again, we have seen that in practice, even at modest permutation order $n$, this class contains couplings that can substantially outperform both i.i.d. walkers and antithetic termination [Reid et al., 2024c].
We now replace $\Lambda_{2}(\eta_{\textrm{G}})$ in Eq. 53 by the set of measures $\{\mu^{(\sigma)}\}_{\sigma\in S_{n}}$ obtained by pushing permutation densities through the required inverse CDF. $S_{n}$ is the set of permutations $[\![n]\!]\to[\![n]\!]$. Denote by $\widehat{\psi}(\cdot):[\![n]\!]\to\mathbb{R}^{N}$ the function
$\widehat{\psi}(q^{(i)})\coloneqq\mathbb{E}_{u\sim\mathcal{U}((\frac{q-1}{n},\frac{q}{n}])}\left(\widetilde{\psi}(F^{-1}_{\eta_{\textrm{G}}}(u)^{(i)})\right).$ (54)
@@
$\sigma^{*}=\arg\min_{\sigma\in S_{n}}\sum_{q_{1}\in[\![n]\!]}\sum_{q_{2}\in[\![n]\!]}\left[\left(\widehat{\psi}(q_{1}^{(i)})+\widehat{\psi}(\sigma(q_{1})^{(i)})\right)^{\top}\left(\widehat{\psi}(q_{2}^{(j)})+\widehat{\psi}(\sigma(q_{2})^{(j)})\right)\right]^{2}$ (55)
@@
Solving the matching problem. Though we can now efficiently estimate the cost function with MC, it is is still difficult to solve the matching problem in Eq. 55 efficiently because it is quadratic [Finke et al., 1987, Burkard et al., 1998]. Minimum weights quadratic bipartite matching problems are generally NP-hard, but here the problem has extra symmetry that makes it simpler.
In App. C.3.1, we discuss possible approaches to directly solve the matching problem in Eq. 55. In particular, we prove that it can be solved with high probability at a time complexity independent of the number of nodes $N$, and give a polynomial time algorithm for the special case of diagonal kernel entries $k(v_{i},v_{i})$. Notwithstanding this progress, as a simpler first recourse we can just make a further approximation by restricting the sum to $q_{1}=q_{2}$ diagonal terms. This is the approximation used to get the results in the main text. Doing so, we arrive at
$\sigma^{*}=\arg\min_{\sigma\in S_{n}}\sum_{q\in[\![n]\!]}\left[\left(\widehat{\psi}(q^{(i)})+\widehat{\psi}(\sigma(q)^{(i)})\right)^{\top}\left(\widehat{\psi}(q^{(j)})+\widehat{\psi}(\sigma(q)^{(j)})\right)\right]^{2}.$ (56)
@@
C.3.1 Proof of Thm. 4.1: towards an analytic solution to Eq. 55
To assuage any possible dissatisfaction with the $q_{1}=q_{2}$ approximation in Eq. 56, in this section we discuss progress towards solving the optimisation problem in Eq. 55 exactly. In particular, we prove Thm. 4.1: that it can be solved with high probability and
@@
    at a time complexity independent from the dimensionality of vectors $\{\widehat{\psi}(q^{(i,j)})\}_{q=1}^{n}$ (at the cost of a one-time pre-processing), or
@@
Proof. First, for clarity of exposition, we will introduce more convenient notation and important definitions. For a given $\epsilon>0$ and a vector $\mathbf{v}\in\mathbb{R}^{d}$, define the set $N_{\epsilon}(\mathbf{v})=\{\mathbf{w}\in\mathbb{R}^{d}:|\alpha_{\mathbf{v},\mathbf{w}}-\frac{\pi}{2}|\leq\epsilon\}$, where $\alpha_{\mathbf{v},\mathbf{w}}$ is an angle between $\mathbf{v}$ and $\mathbf{w}$. Then define the property of $\epsilon$-separation as follows.
@@
Having considered both claims, the proof of Thm. 4.1 is complete. ∎